What is the range of f(x)=-(x+2)^2+8 in set and interval notations
Answer by stanbon(75887)

What is the range of f(x)= -(x+2)^2+8 in set and interval notations
-------
f(x) = -(x^2 + 4x + 4) + 8
-------
f(x) = -x^2 -4x +4
--------------------------
Vertex occurs when x = -b/(2a) = 4/(2*-1) = -2
Maximum y value = f(-2) = -(-2+2)^2+8 = 8
------------
Range = (-oo,8]
--------
Cheers,
Stan H.
------------